OpenStudy (jaylelile):

please help, I do not understand how to do these problems... solve for t. A=P+prt

OpenStudy (caozeyuan):

Is this all you have, what are the symbols stand for?

OpenStudy (jaylelile):

this is all I have... all I understand is that I'm supposed to isolate the t and all the letters will still be there but I don't know I'm still confused.

OpenStudy (caozeyuan):

fist step: prt=A-P

OpenStudy (caozeyuan):

then t=(A-P)/pr

OpenStudy (jaylelile):

ohhhh I see I was only subtracting I didn't divide at all. thank you very much.

OpenStudy (tkhunny):

Generally, it is bad form to use "P" and "p" to mean the same thing. Very confusion. I = P + Prt or I = p + prt Not a mixture.